# LLM-Neo: Parameter Efficient Knowledge

### v1

@@ IV-B Scaling law in LLM-Neo
IV-C Applied on Minitron
We further perform distillation from Llama 3.1 to the Nvidia Minitron 4B [25, 26] using the 50M tokens. All the experiments are conducted on 6 A100 40G GPUs. Fig. 5 shows the results on several benchmarks. We can find that our proposed LLM-Neo performs well on the pruned Minitron 4B, demonstrating its robustness. Specifically, LLM-Neo gets an average score of 54.37, which is 0.52 higher than the base model. The trained weight has been available at this repository.
TABLE II: Memory usage and training time comparison across methods for the Minitron-4B-Width-Base model. Results show LLM-Neo compatibility with existing efficient methods.
@@ IV-C Applied on Minitron
IV-D Compatibility with More Memory Optimizations
To test the robustness of LLM-Neo combined with other memory optimizations, we explore the compatibility with existing LLM optimization techniques when distilling the Minitron 4B. Specifically, we further conduct LLM-Neo with ZeRO-1 methods. As shown in Table II, we can find that LLM-Neo works well with ZeRO1 and ZeRO2, highlighting the relationship between different ZeRO levels and their impact on performance metrics such as memory consumption and time efficiency. Specifically, KD would be out-of-memory though we decrease the batch size to 1.
V Conclusion
In this work, we propose a novel LLM-Neo framework, aiming to efficiently transfer knowledge from a large language model (LLM) teacher to a compact student. We first revisit the knowledge distillation (KD) and low-rank adaption (LoRA), and find that they share the same paradigm. Therefore, we explore the strategy combining LoRA and KD to enhance the efficiency of knowledge transfer. We first summarize some guidelines and further develop the LLM-Neo, i.e., the larger rank around 128 is more suitable. Experimental results on compressing Llama 2 and Llama 3.1 show that LLM-Neo outperforms various baselines. For future work, we would like to explore the performance of LLM-Neo on more training samples and more LoRA variants.

### v2

@@ Appendix B Results on Minitron-4B
Appendix C Compatibility with More Memory Optimizations
To test the robustness of Llm-Neo combined with other memory optimizations, we explore the compatibility with existing LLM optimization techniques when distilling the Minitron 4B. Specifically, we further conduct Llm-Neo with ZeRO-1 methods. As shown in Table 4, we can find that Llm-Neo works well with ZeRO1 and ZeRO2, highlighting the relationship between different ZeRO levels and their impact on performance metrics such as memory consumption and time efficiency. Specifically, KD would be out-of-memory though we decrease the batch size to 1.
@@ Appendix C Compatibility with More Memory Optimizations
Table 4: Memory usage and training time comparison across methods for the Minitron-4B-Width-Base model. Results show LLM-Neo compatibility with existing efficient methods.